OpenStudy (anonymous):

Solve the eqation: a 2 1 ______ + ______ = _____ a^2 -36 a - 6 a + 6

OpenStudy (anonymous):

for starters \[(a^2-36) = (a-6)(a+6)\] then you can find a common denominator and solve for a

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

Could you guys tell me how you got your answer please??

OpenStudy (anonymous):

a should be -9 i think

OpenStudy (anonymous):

you have a/[(a-6)(a+6)] + 2(a+6)/[(a-6)(a+6)] = 1(a-6)/[(a-6)(a+6)] withe [(a-6)(a+6)] as your common denominator

OpenStudy (anonymous):

this becomes (a + 2a + 12)/[(a-6)(a+6)] = (a-6)/[(a-6)(a+6)]

OpenStudy (anonymous):

then (3a+12)/[(a-6)(a+6)] = (a-6)/[(a-6)(a+6)] becomes just (3a+12) = (a-6) by multiplying both sides by the denominator and then you can solve for a

OpenStudy (anonymous):

sure a 2 1 ______ + ______ = _____ a^2 -36 a - 6 a + 6 a ------- + 2/a-6==1/a+6 (a−6)(a+6) taking 1/a-6 common on lhs it becomes (1/a-6)(a/a+6 +2)=1/a+6 (1/a-6)(3a+6/a+6)=1/a+6 cancel a+6 on both sides solve x

OpenStudy (anonymous):

(3a+6)/(a-6)=1 3a+6=a-6 2a=-12 a=-6

OpenStudy (anonymous):

doesnt that make the denominator zero for the rhs of the original equation?

OpenStudy (anonymous):

I think he's right cause it can't be -9. When i put -9 my teacher marked it wrong so it can't be that. But anyways thanks guys

OpenStudy (anonymous):

~.~ im not convinced that this is -6. If you plug in -9 into the original equation you get the right answer http://www.wolframalpha.com/input/?i=%28-9%29%2F%28%28-9%29^2-36%29%2B2%2F%28%28-9%29-6%29+-+1%2F%28%28-9%29%2B6%29 should equal zero when i subtract both sides by 1/(a+6) where a=-9 if a=-6 then you have a denominator being zero which dividing by zero is impossible. if you plug in -9 into a calculator you get that both sides of the equation are the same and -6 is uncomputable since its dividing by zero

OpenStudy (anonymous):

i agree with you

OpenStudy (anonymous):

yea.. i see what your saying